# Union of Finitely Generated Congruences on Ground Term Algebra

### v1

@@ 4. $\mathop{\leftrightarrow}^{*}_{E\cup F}$ has three congruence classes:
{£$\{\,\mathsterling,\;\flat,\;\newline \,f(\#,\#),\;f(\#,\$),\;f(\#,\mathsterling),\;f(\#,\flat),\;\newline f(\$,\#),\;f(\$,\$),\;f(\$,\mathsterling),\;f(\$,\flat),\;\newline f(\mathsterling,\#),\;f(\mathsterling,\$),\;f(\mathsterling,\flat),f(\mathsterling,\mathsterling),\;\newline f(\flat,\#),\;f(\flat,\$),\;f(\flat,\mathsterling),\;f(\flat,\flat)\,\},\newline \{\,\#\,\},\{\,\$\,\}$.
@@
#$\#/_{\Theta\langle E,F\rangle}\subset\#/_{\Theta\langle{E\cup F},\emptyset\rangle}$ and #$\#/_{\Theta\langle{F},E\rangle}\subset\#/_{\Theta\langle{E\cup F},\emptyset\rangle}$.
@@
Let $\Sigma=\Sigma_{0}\cup\Sigma_{2}$, $\Sigma_{0}=\{\,\#,\;\$\,\}$, $\Sigma_{2}=\{\,f\,\}$,
$E=\{\,f(\#,\#)\doteq\$,\;f(\#,\$)\doteq\$,\;f(\$,\#)\doteq\$,\;f(\$,\$)\doteq\$\,\}$ and $F=\{\,f(\#,\#)\doteq\#\,\}$.
@@
$ST\langle E\rangle=\{\,\#,\;\$,\;f(\#,\#),\;f(\#,\$),\;f(\$,\#),\;f(\$,\$)\,\}$,  $ST\langle F\rangle=\{\,\#,\;f(\#,\#)\,\}$,  $W=ST\langle E\rangle$.
@@
The equivalence classes of $\Theta\langle F,E\rangle$ are {#$\{\,\#,\;f(\#,\#)\,\},\;\{\,\$\,\},\;\{\,f(\#,\$)\,\},\;\{\,f(\$,\#)\,\},\;\{\,f(\$,\$)\,\}$.
The only equivalence class of $\Theta\langle E\cup F,\emptyset\rangle$ is {#,\$$\{\,\#,\;\$,\;f(\#,\#),\;f(\#,\$),\;f(\$,\#),\;f(\$,\$)\,\}$.
Observe that $E\not\subseteq\mathop{\leftrightarrow}^{*}_{F}$ and $F\not\subseteq\mathop{\leftrightarrow}^{*}_{E}$. Since (#,\$$(\#,\$)\not\in\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$, and (#,\$$(\#,\$)\in\mathop{\leftrightarrow}^{*}_{E\cup F}$, we have $\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}\subset\mathop{\leftrightarrow}^{*}_{E\cup F}$.
@@
Let $\Sigma=\Sigma_{0}\cup\Sigma_{1}\cup\Sigma_{2}$, $\Sigma_{0}=\{\,\#,\;\$\,\}$, $\Sigma_{1}=\{\,g\,\}$, $\Sigma_{2}=\{\,f\,\}$,
$E=\{\,f(\#,\#)\doteq g(\#),\;f(\#,g(\#))\doteq g(\#),\;f(g(\#),\#)\doteq g(\#),\;f(g(\#),g(\#))\doteq g(\#),\;g(g(\#))\doteq g(\#),\;\newline f(\#,\$)\doteq\$,\;f(\$,\#)\doteq\$,\;f(\$,\$)\doteq\$,\;g(\$)\doteq\$\,\}$.
@@
Intuitively, $F$ can simulate the first five equations of $E$, that is, the congruence generated by the first five equations of $E$ is a proper subset of $\mathop{\leftrightarrow}^{*}_{F}$. We need the constant \$currency-dollar\\$\$ and the last four equations of $E$ to ensure that $\mathop{\leftrightarrow}^{*}_{E}\not\subseteq\mathop{\leftrightarrow}^{*}_{F}$. Note the following equations:
  • $ST\langle E\rangle=\{\,\#,\;\$,\;f(\#,\#),\;f(\#,\$),\;f(\$,\#),\;f(\$,\$),\;g(\#),\;g(\$),\;f(\#,g(\#)),\;f(g(\#),\#),\;f(g(\#),g(\#)),\;g(g(\#))\,\}$,
@@
\$$\$/_{\mathop{\leftrightarrow}^{*}_{E}}=\$/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}=\{\,t\in T_{\Sigma}\mid\mbox{ there exists }\alpha\in POS(t)\mbox{ such that }t|_{\alpha}=\$\,\}$.
#$\#/_{\mathop{\leftrightarrow}^{*}_{F}}=\#/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}=\{\,t\in T_{\Sigma}\mid\mbox{ there does not exist }\alpha\in POS(t)\mbox{ such that }t|_{\alpha}=\$\,\}$.
@@
Let $\Sigma=\Sigma_{0}\cup\Sigma_{2}$, $\Sigma_{0}=\{\,\#,\;\$\,\}$, $\Sigma_{2}=\{\,f\,\}$,
$E=\{\,f(\#,\#)\doteq f(\#,\$),\;f(\#,\#)\doteq f(\$,\#),\;f(\#,\#)\doteq f(\$,\$)\,\}$ and $F=\{\,\#\doteq\$\,\}$.
@@
$ST\langle E\rangle=\{\,\#,\;\$,\;f(\#,\#),\;f(\#,\$),\;f(\$,\#),\;f(\$,\$)\,\}$,  $ST\langle F\rangle=\{\,\#,\$\,\}$,   $W=ST\langle E\rangle$,  $E\subseteq\mathop{\leftrightarrow}^{*}_{F}$,   $\mathop{\leftrightarrow}^{*}_{E\cup F}=\mathop{\leftrightarrow}^{*}_{F}$.
@@
The equivalence classes of $\Theta\langle E,F\rangle$ are {#},{\$},$\{\,\#\,\},\;\{\,\$\,\},\;\{\,f(\#,\#),\;f(\#,\$),\;f(\$,\#),\;f(\$,\$)\,\}$.
The equivalence classes of $\Theta\langle F,E\rangle$ are {#,\$},$\{\,\#,\;\$\,\},\;\{\,f(\#,\#),\;f(\#,\$),\;f(\$,\#),\;f(\$,\$)\,\}$.
The equivalence classes of $\Theta\langle E\cup F,\emptyset\rangle$ are {#,\$},$\{\,\#,\;\$\,\},\;\{\,f(\#,\#),\;f(\#,\$),\;f(\$,\#),\;f(\$,\$)\,\}$.
@@
Let $\Sigma=\Sigma_{0}\cup\Sigma_{2}$, $\Sigma_{0}=\{\,\#,\;\flat,\;\$,\;\mathsterling\,\}$, $\Sigma_{2}=\{\,f\,\}$,
$E=\{\,\#\doteq\$\,\}$, and $F=\{\,\mathsterling\doteq\flat\,\}$.
We remark that $f(\#,\mathsterling)\mathop{\leftrightarrow}^{*}_{E\cup F}f(\$,\flat)$,  $(f(\#,\mathsterling),f(\$,\flat))\not\in\mathop{\leftrightarrow}^{*}_{E}$,   and $(f(\#,\mathsterling),f(\$,\flat))\not\in\mathop{\leftrightarrow}^{*}_{F}$. Thence
@@
$ST\langle E\rangle=\{\,\#,\;\$\,\}$,   $ST\langle F\rangle=\{\,\mathsterling,\;\flat\,\}$,   $W=\{\,\#,\;\$,\;\mathsterling,\;\flat\,\}$.
@@ 4 Fast Ground Completion Introducing Additional Constants
From now on throughout this section, let $\Sigma$ be a signature, $E$ and $F$ be GTESs over $\Sigma$, $W=ST(E\cup F)$, and $n=size(E)+size(F)$. We recall the fast ground completion algorithm of Fülöp and Vágvölgyi [2]. Its input is $\Sigma$, $E$, and $F$. Their fast ground completion algorithm [2] produces in $O(n\,\mathrm{log}\;n)$ time a finite set of new constants and a reduced GTRS $R\langle E,F\rangle$ over $\Sigma$ extended with the new constants such that $\mathop{\leftrightarrow}^{*}_{E}=\mathop{\leftrightarrow}^{*}_{R\langle E,F\rangle}\cap T_{\Sigma}\times T_{\Sigma}$. Their algorithm calls a congruence closure algorithm on the subterm dag for $E\cup F$ [4], and runs in $O(n\,\mathrm{log}\,n)$ time by Proposition 2.16.
We now cite the Claim on [17, p. 167].
Proposition 4.25.
[17] $\Theta\langle E,F\rangle=\{\,(\widehat{x},\widehat{y})\mid(x,y)\in\rho\langle E,F\rangle\,\}$.
@@
By Proposition 4.25, for each equivalence class $\vartheta\in W/_{\Theta\langle E,F\rangle}$, there exists exactly one equivalence class $\eta\in V/_{\rho\langle E,F\rangle}$ such that $\vartheta=\{\,\widehat{x}\mid x\in\eta\,\}$. To simplify our notations, we give $\vartheta$ the unique name of $\eta$. The elements of $C\langle E,F\rangle$ are viewed as constants, and we define the signature $\Sigma\cup C\langle E,F\rangle$.
Statement 4.27.
$|C\langle E,F\rangle|\leq|W|\leq size(W)\leq n$.
Proof.
By Definition 4.26. ∎
Definition 4.28.
[2] Let
@@
be a GTRS over the ranked alphabet $\Sigma\cup C\langle E,F\rangle$.
@@
We say that GTRS $R\langle E,F\rangle$ is total if for any $\sigma\in\Sigma_{m}$ with $m\in{\mathbb{N}}$, and $a_{1},\ldots,a_{m}\in C\langle E,F\rangle$, there exists a rewrite rule in GTRS $R\langle E,F\rangle$ with left-hand side $\sigma(a_{1},\ldots,a_{m})$.
@@
Proposition 4.30.
[2] $R\langle E,F\rangle$ is reduced.
@@
Proposition 4.32.
[2] For every $t\in W$, we have $t\mathop{\rightarrow}_{R\langle E,F\rangle}^{*}t/_{\Theta\langle E,F\rangle}$.
@@
Let $t\in W$. By Proposition 4.32, we have $t\mathop{\rightarrow}_{R\langle E,F\rangle}^{*}t/_{\Theta\langle E,F\rangle}$. By direct inspection of the rewrite rules of $R\langle E,F\rangle$, we get that for each $t\in W$, $t/_{\Theta\langle E,F\rangle}$ is irreducible for $R\langle E,F\rangle$. Accordingly $t/_{\Theta\langle E,F\rangle}$ is a $\mathop{\rightarrow}_{R\langle E,F\rangle}$ normal form of $t$. By Proposition 2.3, Proposition 4.31 $t\in W$ has a unique normal form, $t{\downarrow}_{R\langle E,F\rangle}$. Then we have $t{\downarrow}_{R\langle E,F\rangle}=t/_{\Theta\langle E,F\rangle}$. ∎
@@
[2] For all $s\in T_{\Sigma}$ and $t\in W$,
Proposition 4.35.
[2] For all $s,t\in T_{\Sigma}$,
Proposition 4.36.
[2] For all $s\in T_{\Sigma}$ and $t\in W$,
@@
Let $\sigma\in\Sigma_{m}$ with $m\in{\mathbb{N}}$, and let $t_{1}/_{\Theta\langle E,F\rangle},\ldots,t_{m}/_{\Theta\langle E,F\rangle}$ be equivalence classes of $\Theta\langle E,\;F\rangle$ for some $t_{1},\ldots t_{m}\in W$. By our assumption, there exists $s\in W$ such that $\sigma(t_{1},\ldots,t_{m})\mathop{\leftrightarrow}^{*}_{E}s$. By Proposition 4.35, $\sigma(t_{1},\ldots,t_{m}){\downarrow}_{R\langle E,F\rangle}=s{\downarrow}_{R\langle E,F\rangle}$. By Proposition 4.33 $s{\downarrow}_{R\langle E,F\rangle}=s/_{\Theta\langle E,F\rangle}$. Thus $\sigma(t_{1},\ldots,t_{m}){\downarrow}_{R\langle E,F\rangle}=s/_{\Theta\langle E,F\rangle}$. Therefore,
$\sigma(t_{1},\ldots,t_{m})\mathop{\rightarrow}_{R\langle E,F\rangle}^{*}\sigma(t_{1}{\downarrow}_{R\langle E,F\rangle},\ldots,t_{m}{\downarrow}_{R\langle E,F\rangle})=$ (by Proposistion 4.32)
@@
Base case: Let $height(s)=0$. Then $s=\sigma$ for some $\sigma\in\Sigma_{0}$. By our assumption that $R\langle E,F\rangle$ is a total GTES and by Definitions 4.28 and 4.29, there is a rewrite rule $\sigma\dot{\rightarrow}\sigma/_{\Theta\langle E,F\rangle}$ in $R\langle E,F\rangle$. Then $\sigma\in W$ and $s{\downarrow}_{R\langle E,F\rangle}=\sigma/_{\Theta\langle E,F\rangle}$.
[…]
@@
[…]
@@
By direct inspection of Definition 4.28, we get that $p_{\Theta\langle E,F\rangle}$ is irreducible for $R\langle E,F\rangle$. Thus $s{\downarrow}_{R\langle E,F\rangle}=\sigma(p_{1},\ldots,p_{m})/_{\Theta\langle E,F\rangle}$. ∎
@@
It is a consequence of Lemma 4.40 and Lemma 4.41 and Proposition 4.35. ∎
Statement 4.43.
GTRS $R\langle E,F\rangle$ reach a constant $a\in C\langle E,F\rangle$ from a ground term $t$ over the ranked alphabet $\Sigma\cup C\langle E,F\rangle$ if and only if $t{\downarrow}_{R\langle E,F\rangle}=a$.
Proof.
()($\Rightarrow$)( ⇒ ) The constant $a\in C\langle E,F\rangle$ is irreducibe for $R\langle E,F\rangle$ by direct inspection of Definition 4.28. By Proposition 4.31, ${R\langle E,F\rangle}$ is convergent. Accordingly $a\in C\langle E,F\rangle$ is the unique $R\langle E,F\rangle$-normal form of $t$.
()($\Leftarrow$)( ⇐ ) This direction is trivial. ∎
@@
$BSTEP\langle E,F\rangle=\{\,(a,b)\in C\langle E,F\rangle\times C\langle E,F\rangle\mid\mbox{ there exists a rewrite rule }\sigma(a_{1},\ldots,a_{m})\dot{\rightarrow}a\in R\langle E,F\rangle\mbox{ and }1\leq i\leq m\mbox{ such that }b=a_{i}\,\}$.
@@
Statement 4.45.
Let $n=size(E)+size(F)$. Then
1. $|R\langle E,F\rangle|\leq|W|\leq size(W)\leq size(E\cup F)\leq size(E)+size(F)=n$.
2. $size(R\langle E,F\rangle)\leq size(W)+|W|\leq 2\cdot n$  and
3. $|BSTEP\langle E,F\rangle|\leq size(R\langle E,F\rangle)\leq 2\cdot n$.
@@
By direct inspection of Definition 4.28 and by Statement 2.7 we have Statement 1.
[…]
By direct inspection of Definition 4.44 and by Statement 2 we have Statement 3.
@@
Definition 4.46.
Algorithm Fast Ground Completion Introducing Additional Constants (FGC for short) [2]

### v2

@@ 4. $\mathop{\leftrightarrow}^{*}_{E\cup F}$ has three congruence classes:
$\{\,\mathsterling,\;\flat,\;\newline \,f(\#,\#),\;f(\#,\mathdollar),\;f(\#,\mathsterling),\;f(\#,\flat),\;\newline f(\mathdollar,\#),\;f(\mathdollar,\mathdollar),\;f(\mathdollar,\mathsterling),\;f(\mathdollar,\flat),\;\newline f(\mathsterling,\#),\;f(\mathsterling,\mathdollar),\;f(\mathsterling,\flat),f(\mathsterling,\mathsterling),\;\newline f(\flat,\#),\;f(\flat,\mathdollar),\;f(\flat,\mathsterling),\;f(\flat,\flat)\,\},\newline \{\,\#\,\},\{\,\mathdollar\,\}$.
@@
$_{\Theta\langle E,F\rangle}\subset\#/_{\Theta\langle{E\cup F},\emptyset\rangle}$ and $_{\Theta\langle{F},E\rangle}\subset\#/_{\Theta\langle{E\cup F},\emptyset\rangle}$.
@@
Let $\Sigma=\Sigma_{0}\cup\Sigma_{2}$, $\Sigma_{0}=\{\,\#,\;\mathdollar\,\}$, $\Sigma_{2}=\{\,f\,\}$,
[…]
@@
[…]
Observe that $E\not\subseteq\mathop{\leftrightarrow}^{*}_{F}$ and $F\not\subseteq\mathop{\leftrightarrow}^{*}_{E}$. Since $(\#,\mathdollar)\not\in\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}$, and $(\#,\mathdollar)\in\mathop{\leftrightarrow}^{*}_{E\cup F}$, we have $\mathop{\leftrightarrow}^{*}_{E}\cup\mathop{\leftrightarrow}^{*}_{F}\subset\mathop{\leftrightarrow}^{*}_{E\cup F}$.
@@
[…]
@@
[…]
  • 1.
    $ST\langle E\rangle=\{\,\#,\;\mathdollar,\;f(\#,\#),\;f(\#,\mathdollar),\;f(\mathdollar,\#),\;f(\mathdollar,\mathdollar),\;g(\#),\;g(\mathdollar),\;f(\#,g(\#)),\;f(g(\#),\#),\;f(g(\#),g(\#)),\;g(g(\#))\,\}$,
  • 2.
@@
$\mathdollar/_{\mathop{\leftrightarrow}^{*}_{E}}=\mathdollar/_{\mathop{\leftrightarrow}^{*}_{E\cup F}}=\{\,t\in T_{\Sigma}\mid\mbox{ there exists }\alpha\in POS(t)\mbox{ such that }t|_{\alpha}=\mathdollar\,\}$.
[…]
@@
Let $\Sigma=\Sigma_{0}\cup\Sigma_{2}$, $\Sigma_{0}=\{\,\#,\;\mathdollar\,\}$, $\Sigma_{2}=\{\,f\,\}$,
[…]
@@
[…]
@@
[…]
@@
[…]
$E=\{\,\#\doteq\mathdollar\,\}$, and $F=\{\,\mathsterling\doteq\flat\,\}$.
[…]
@@
[…]
@@ 4 Fast Ground Completion Introducing Additional Constants
From now on throughout the paper, let $\Sigma$ be a signature, $E$ and $F$ be GTESs over $\Sigma$, $W=ST\langle E\cup F\rangle$, and $n=\mathrm{size}(E)+\mathrm{size}(F)$. We recall the fast ground completion algorithm of Fülöp and Vágvölgyi [5, Algorithm 3.6.]. Its input is $\Sigma$, $E$, and $F$. Their fast ground completion algorithm [5, Algorithm 3.6.] produces in $O(n\,\mathrm{log}\,n)$ time a finite set of new constants and a reduced GTRS $R\langle E,F\rangle$ over $\Sigma$ extended with the new constants such that $\mathop{\leftrightarrow}^{*}_{E}=\mathop{\leftrightarrow}^{*}_{R\langle E,F\rangle}\cap T_{\Sigma}\times T_{\Sigma}$. Their algorithm calls a congruence closure algorithm on the subterm dag for $E\cup F$ [9], and runs in $O(n\,\mathrm{log}\,n)$ time by Proposition 2.17.
We now cite the Claim on [1, p. 167].
Proposition 4.26 ([1, Claim on page 167]).
$\Theta\langle E,F\rangle=\{\,(\widehat{x},\widehat{y})\mid(x,y)\in\rho\langle E,F\rangle\,\}$.
@@
By Proposition 4.26, for each equivalence class $\vartheta\in W/_{\Theta\langle E,F\rangle}$, there exists exactly one equivalence class $\eta\in V/_{\rho\langle E,F\rangle}$ such that $\vartheta=\{\,\widehat{x}\mid x\in\eta\,\}$. To simplify our notations, we give $\vartheta$ the unique name of $\eta$. The elements of $C\langle E,F\rangle$ are viewed as constants, and we define the signature $\Sigma\cup C\langle E,F\rangle$. By Definition 4.27, we have the following statement.
Statement 4.28.
$|C\langle E,F\rangle|\leq|W|\leq\mathrm{size}(W)\leq n$.
Definition 4.29 ([5, Section 3]).
Let
@@
be a GTRS over the signature $\Sigma\cup C\langle E,F\rangle$.
@@
We say that GTRS $R\langle E,F\rangle$ is total if for any $\sigma\in\Sigma_{m}$ with $m\in{\mathbb{N}}$, and ${b}_{1},\ldots,{b}_{m}\in C\langle E,F\rangle$, there exists a rewrite rule in GTRS $R\langle E,F\rangle$ with left-hand side $\sigma({b}_{1},\ldots,{b}_{m})$.
@@
Proposition 4.31 ([5, Proposition 3.1]).
$R\langle E,F\rangle$ is reduced.
@@
Proposition 4.33 ([5, Proposition 3.2]).
For every $t\in W$, we have $t\mathop{\rightarrow}_{R\langle E,F\rangle}^{*}t/_{\Theta\langle E,F\rangle}$.
@@
Let $t\in W$. By Proposition 4.33, we have $t\mathop{\rightarrow}_{R\langle E,F\rangle}^{*}t/_{\Theta\langle E,F\rangle}$. By direct inspection of the rewrite rules of $R\langle E,F\rangle$, we get that for each $t\in W$, $t/_{\Theta\langle E,F\rangle}$ is irreducible for $R\langle E,F\rangle$. Accordingly $t/_{\Theta\langle E,F\rangle}$ is a $\mathop{\rightarrow}_{R\langle E,F\rangle}$ normal form of $t$. By Proposition 2.4 and Proposition 4.32, $t\in W$ has a unique normal form, $t{\downarrow}_{R\langle E,F\rangle}$. Then we have $t{\downarrow}_{R\langle E,F\rangle}=t/_{\Theta\langle E,F\rangle}$. ∎
Proposition 4.35 ([5, Lemma 3.3]).
For all $s\in T_{\Sigma}$ and $t\in W$,
Proposition 4.36 ([5, Theorem 3.4]).
Proposition 4.37 ([5, Theorem 3.5]).
For all $s,t\in T_{\Sigma}$,
Proposition 4.34 and Proposition 4.37 imply the following proposition.
@@
For all $s\in T_{\Sigma}$ and $t\in W$,
@@
Let $\sigma\in\Sigma_{m}$ with $m\in{\mathbb{N}}$, and let $t_{1}/_{\Theta\langle E,F\rangle},\ldots,t_{m}/_{\Theta\langle E,F\rangle}$ be equivalence classes of $\Theta\langle E,\;F\rangle$ for some $t_{1},\ldots t_{m}\in W$. By our assumption, there exists $s\in W$ such that $\sigma(t_{1},\ldots,t_{m})\mathop{\leftrightarrow}^{*}_{E}s$. By Proposition 4.37, $\sigma(t_{1},\ldots,t_{m}){\downarrow}_{R\langle E,F\rangle}=s{\downarrow}_{R\langle E,F\rangle}$. By Proposition 4.34, $s{\downarrow}_{R\langle E,F\rangle}=s/_{\Theta\langle E,F\rangle}$. Thus $\sigma(t_{1},\ldots,t_{m}){\downarrow}_{R\langle E,F\rangle}=s/_{\Theta\langle E,F\rangle}$. Therefore,
$\sigma(t_{1},\ldots,t_{m})\mathop{\rightarrow}_{R\langle E,F\rangle}^{*}\sigma(t_{1}{\downarrow}_{R\langle E,F\rangle},\ldots,t_{m}{\downarrow}_{R\langle E,F\rangle})=$ (by Proposition 4.33)
@@
Base Case: Let $height(s)=0$. Then $s=\sigma$ for some $\sigma\in\Sigma_{0}$. By our assumption that $R\langle E,F\rangle$ is a total GTES and by Definitions 4.29 and 4.30, there is a rewrite rule $\sigma\dot{\rightarrow}\sigma/_{\Theta\langle E,F\rangle}$ in $R\langle E,F\rangle$. Then $\sigma\in W$ and $s{\downarrow}_{R\langle E,F\rangle}=\sigma/_{\Theta\langle E,F\rangle}$.
Induction Step: Let $k\in{\mathbb{N}}$, and let us assume that we have shown the lemma for all integers less than or equal to $k$. Let $height(s)=k+1$. Then $s=\sigma(s_{1},\ldots,s_{m})$ for some $\sigma\in\Sigma_{m}$, $m\geq 1$, and $s_{1},\ldots,s_{m}\in T_{\Sigma}$. By the induction hypothesis, for each $i\in\{\,1,\ldots,m\,\}$, there exist $t_{i}\in W$ such that $s_{i}{\downarrow}_{R\langle E,F\rangle}=t_{i}/_{\Theta\langle E,F\rangle}$. By our assumption that $R\langle E,F\rangle$ is a total GTRS, there is a rewrite rule with left-hand side $\sigma(t_{1}/_{\Theta\langle E,F\rangle},\ldots,t_{m}/_{\Theta\langle E,F\rangle})$ in $R\langle E,F\rangle$. By Definition 4.29, this rule is of the form
@@
in $R\langle E,F\rangle$ for some $p_{i}\in t_{i}/_{\Theta\langle E,F\rangle}$ for every $i\in\{\,1,\ldots,m\,\}$. Moreover $\sigma(p_{1},\ldots,p_{m})\in W$. Then
@@
By direct inspection of Definition 4.29, we get that $\sigma(p_{1},\ldots,p_{m})/_{\Theta\langle E,F\rangle}$ is irreducible for $R\langle E,F\rangle$. Thus $s{\downarrow}_{R\langle E,F\rangle}=\sigma(p_{1},\ldots,p_{m})/_{\Theta\langle E,F\rangle}$. ∎
The following proposition is a consequence of Lemma 4.41 and Lemma 4.42 and Proposition 4.37.
@@
Statement 4.44.
GTRS $R\langle E,F\rangle$ reaches a constant ${b}\in C\langle E,F\rangle$ from a ground term $t$ over the signature $\Sigma\cup C\langle E,F\rangle$ if and only if $t{\downarrow}_{R\langle E,F\rangle}={b}$.
@@
$(\Rightarrow)$ The constant ${b}\in C\langle E,F\rangle$ is irreducibe for $R\langle E,F\rangle$ by direct inspection of Definition 4.29. By Proposition 4.32, $R\langle E,F\rangle$} is convergent. Accordingly ${b}\in C\langle E,F\rangle$ is the unique $R\langle E,F\rangle$-normal form of $t$.
$(\Leftarrow)$ This direction is trivial. ∎
@@
[…]
@@
Statement 4.46.
Let $n=\mathrm{size}(E)+\mathrm{size}(F)$. Then
1. $|R\langle E,F\rangle|\leq|W|\leq\mathrm{size}(W)\leq size(E\cup F)\leq\mathrm{size}(E)+\mathrm{size}(F)=n$.
2. $\mathrm{size}(R\langle E,F\rangle)\leq\mathrm{size}(W)+|W|\leq 2n$  and
3. $|BSTEP\langle E,F\rangle|\leq\mathrm{size}(R\langle E,F\rangle)\leq 2n$.
@@
By direct inspection of Definition 4.29 and by Statement 2.8 we have Statement 1.
We now show Statement 2. By Definition 4.29, for each rule $\sigma({b}_{1},\ldots,{b}_{m})\dot{\rightarrow}{b}\in R\langle E,F\rangle$, there exists a ground term $\sigma(t_{1},\ldots,t_{m})\in W$ such that ${b}_{i}=t_{i}/_{\Theta\langle E,F\rangle}$ for every $i\in\{\,1,\ldots,m\,\}$. Then we assign to the occurrence of ${b}_{i}$ the root symbol of the ground term $t_{i}$ for every $i\in\{\,1,\ldots,m\,\}$. Consequently $\mathrm{size}(R\langle E,F\rangle)\leq\mathrm{size}(W)+|W|$. By Statement 1, $\mathrm{size}(W)+|W|\leq 2n$.
By direct inspection of Definition 4.45 and by Statement 2 we have Statement 3. ∎
@@
Definition 4.47 ([5, Algorithm 3.6]).
Algorithm Fast Ground Completion Introducing Additional Constants (FGC for short)